# ProxFly: Robust Control for Close Proximity Quadcopter Flight via Residual Reinforcement Learning

### v1

@@ Abstract
This paper proposes the ProxFly, a residual deep Reinforcement Learning (RL)-based controller for close proximity quadcopter flight. Specifically, we design a residual module on top of a cascaded controller (denoted as basic controller) to generate high-level control commands, which compensate for external disturbances and thrust loss caused by downwash effects from other quadcopters. First, our method takes only the ego state and controllers’ commands as inputs and does not rely on any communication between quadcopters, thereby reducing the bandwidth requirement. Through domain randomization, our method relaxes the requirement for accurate system identification and fine-tuned controller parameters, allowing it to adapt to changing system models. Meanwhile, our method not only reduces the proportion of unexplainable signals from the black box in control commands but also enables the RL training to skip the time-consuming exploration from scratch via guidance from the basic controller. We validate the effectiveness of the residual module in the simulation with different proximities. Moreover, we conduct the real close proximity flight test to compare ProxFly with the basic controller and an advanced model-based controller with complex aerodynamic compensation. Finally, we show that ProxFly can be used for challenging quadcopter in-air docking, where two quadcopters fly in extreme proximity, and strong airflow significantly disrupts flight. However, our method can stabilize the quadcopter in this case and accomplish docking. The resources are available at https://github.com/ruiqizhang99/ProxFly.
@@ I Introduction
Recently, deep learning (DL) is a rapidly developing data-driven method for extracting complex latent representation from large amounts of data[9]. It enables many challenging tasks like visual-based navigation[10, 11] and near-ground flight[12], which provide new perspectives for achieving robust flight control. For example, learning a residual model to compensate for ground effects from real flight data and assist the stable control under complex aerodynamics effect[12]. However, these supervised learning methods rely on extensive data collected from human pilots and manual annotations, and the performance of learned controllers heavily depends on the skill level of the human pilots. DL can also be used to predict the downwash effect through real-world flight datasets, so that we can incorporate the learned model into motion planning, which are verified on high-speed navigation[13], homogeneous[14] and heterogeneous[15] quadcopter swarms. However, these methods are only validated on micro quadcopters with weak downwash effect and large spacing, so it is uncertain whether the conclusions can be generalized to regular quadcopters and other closer proximity flight tasks.
@@ II-B Training Settings
We train two independent full-enveloped flight policies for both quadcopters with the proximal policy optimization (PPO)[29] algorithm. The policy and value networks are three-layer multilayer perceptrons (MLPs) with $128$ units. The first two layers use LeakyReLU[30] as the activation, while the last layer uses the $\tanh$ activation to normalize the output. The Adam optimizer[31] is employed for training these networks. For each episode during the training process, we simulate $20$ seconds of flight (i.e., $10,000$ time steps at the simulation frequency of $500$). In the first $15$ seconds, the takeoff and hover phases are simulated, followed by the landing phase in the last $5$ seconds. At the beginning of each episode, the vehicle’s position is randomly initialized on the ground within a $2m\times 2m$ area, and we set the desired hovering position as $(0,0,1.2)$. During the last 5 seconds, the quadcopter is required to descend vertically at a speed of $0.2$m/s. We train the policy for $500$ epochs with $10$ episodes in each one, which can be finished in $15$ minutes on a laptop with a Intel i7-13700H CPU.
@@ II-B3 Randomization
Propeller Thrust Factor 1.145e-7 7.640e-6
@@
For the proximity flight, the disturbance primarily comes from the downwash flow generated by the rotors. On the one hand, the downwash flow reduces the propeller efficiency of other quadcopters in the wind field, so the actual thrust decreases compared with it in the static air. On the other hand, quadcopter in wind field experiences external forces and torques acting on the body[6]. To simulate the propeller thrust loss and imbalance caused by downwash flow, we randomly initialize a propeller thrust factor from $\mathcal{U}(0.6,1.2)$ for four propellers independently. Meanwhile, although we can calculate the force and torque disturbance with the CFD software or wind tunnel, these methods are resource-intensive, time-consuming, and cannot cover all possible flight scenarios. Therefore, we propose to present the forces in the $x,y,z$ directions as three independent triangular functions with random periods and amplitudes, as shown in Table I. Since downwash flow mainly acts in the vertical direction, the force disturbance in the $z$ direction is more significant. For the torque disturbances around three axes, we simulate them using three independent Gaussian noises. Overall, this simulation approach allows quadcopters to resist airflow disturbances without any communication. In the next section, we conduct the simulation experiment to demonstrate that the policy trained with the proposed noise can resist diverse high-fidelity airflow disturbances.
@@ III-A Simulation Experiments and Analysis
We conduct an experiment in simulator to verify that the proposed disturbances contribute to making the residual module adaptive to high-fidelity simulated disturbances. Based on conclusions from previous research[6], we approximate the effect of real downwash flow as a function of the relative positions and model parameters of two quadcopters. In our setup, the LQ takes off from the ground and hovers at $(0,0,1.2)$, while the SQ takes off vertically from $(-1,0,0)$ and flies along the positive $x$-axis at a speed of $0.2$ m/s at a certain height. After hovering above the LQ for 5 seconds, the SQ continues along the positive $x$-axis and lands vertically at $(1,0,0)$. We set the vertical distance between the SQ and LQ as $\left\{0.25,0.5,0.75\right\}$ to reveal the effect of downwash flow disturbances with different magnitudes and their corresponding compensatory behaviors generated by the residual module.
@@ III-B2 Quadcopter Aerial Docking
Figure 5: The procedure of quadcopter in-air docking. Stage 1: Taking off and hovering. The residual RL controller only assists the basic controller on the large quadcopter (LQ) to achieve faster responses. Stage 2: Small quadcopter approaching. Small quadcopter (SQ) hovers above the LQ and vertically approaches it, which generates strong downwash flow disturbances, and the residual RL controller generates thrust and rate compensation to help stabilize LQ’s position and attitude. (LPF: low-pass filtered) Stage 3: Docked with LQ. The SQ falls freely from 5cm above the LQ, gets docked with the LQ and generates an impulse. The overall mass changes and the thrust compensation from the RL controller on LQ reaches the peak. Stage 4: SQ leaving. SQ takes off from the LQ vertically, and the downwash airflow reappears and gradually decreases while the thrust compensation from the RL controller also gradually decreases.
@@ IV Conclusion
This paper proposes the ProxFly, which leverages a cascaded controller combined with residual RL for diverse close proximity tasks, including hovering, trajectory tracking, in-air docking, and undocking. This approach does not rely on any communication or accurate knowledge of system parameters but only uses a lightweight MLP to learn how to compensate for high-level commands. ProxFly leverages the strengths of both classical control and deep RL. On the one hand, it requires less domain expertise and avoids time-consuming controller parameter tuning. On the other hand, compared to end-to-end RL, it improves sampling efficiency and reduces the proportion of unexplained signals from the black box in the overall control signals.

### v2

@@ Abstract
This paper proposes the ProxFly, a residual deep Reinforcement Learning (RL)-based controller for close proximity quadcopter flight. Specifically, we design a residual module on top of a cascaded controller (denoted as basic controller) to generate high-level control commands, which compensate for external disturbances and thrust loss caused by downwash effects from other quadcopters. First, our method takes only the ego state and controllers’ commands as inputs and does not rely on any communication between quadcopters, thereby reducing the bandwidth requirement. Through domain randomization, our method relaxes the requirement for accurate system identification and fine-tuned controller parameters, allowing it to adapt to changing system models. Meanwhile, our method not only reduces the proportion of unexplainable signals from the black box in control commands but also enables the RL training to skip the time-consuming exploration from scratch via guidance from the basic controller. We validate the effectiveness of the residual module in the simulation with different proximities. Moreover, we conduct the real close proximity flight test to compare ProxFly with the basic controller and an advanced model-based controller with complex aerodynamic compensation. Finally, we show that ProxFly can be used for challenging quadcopter mid-air docking, where two quadcopters fly in extreme proximity, and strong airflow significantly disrupts flight. However, our method can stabilize the quadcopter in this case and accomplish docking. The resources are available at https://github.com/ruiqizhang99/ProxFly.
@@ I Introduction
Recently, deep learning (DL) is a rapidly developing data-driven method for extracting complex latent representation from large amounts of data[9]. It enables many challenging tasks like visual-based navigation[10, 11], close-proximity[12] and near-ground flight[13], which provide new perspectives for achieving robust flight control. For example, learning a residual model to compensate for ground effects from real flight data and assist the stable control under complex aerodynamics effect[13]. However, these supervised learning methods rely on extensive data collected from human pilots and manual annotations, and the performance of learned controllers heavily depends on the skill level of the human pilots. DL can also be used to predict the downwash effect through real-world flight datasets, so that we can incorporate the learned model into motion planning, which are verified on high-speed navigation[14], homogeneous[15] and heterogeneous[16] quadcopter swarms. However, these methods are only validated on micro quadcopters with weak downwash effect and large spacing, so it is uncertain whether the conclusions can be generalized to regular quadcopters and other closer proximity flight tasks.
@@ II-B Training Settings
We train two independent full-enveloped flight policies for both quadcopters with the proximal policy optimization (PPO)[30] algorithm. The policy and value networks are three-layer multilayer perceptrons (MLPs) with $128$ units. The first two layers use LeakyReLU[31] as the activation, while the last layer uses the $\tanh$ activation to normalize the output. The Adam optimizer[32] is employed for training these networks. For each episode during the training process, we simulate $20$ seconds of flight (i.e., $10,000$ time steps at the simulation frequency of $500$). In the first $15$ seconds, the takeoff and hover phases are simulated, followed by the landing phase in the last $5$ seconds. At the beginning of each episode, the vehicle’s position is randomly initialized on the ground within a $2m\times 2m$ area, and we set the desired hovering position as $(0,0,1.2)$. During the last 5 seconds, the quadcopter is required to descend vertically at a speed of $0.2$m/s. We train the policy for $500$ epochs with $10$ episodes in each one, which can be finished in $15$ minutes with parallelized computing [33] on a laptop with a Intel i7-13700H CPU.
@@ II-B3 Randomization
Propeller Thrust Efficiency 1.145e-7 7.640e-6
@@
For the proximity flight, the disturbance primarily comes from the downwash flow generated by the rotors. On the one hand, the downwash flow reduces the propeller efficiency of other quadcopters in the wind field, so the actual thrust decreases compared with it in the static air. On the other hand, quadcopter in wind field experiences external forces and torques acting on the body[6]. To simulate the propeller thrust loss and imbalance caused by downwash flow, we randomly initialize a propeller thrust factor from $\mathcal{U}(0.6,1.2)$ for four propellers independently. Meanwhile, although we can calculate the force and torque disturbance with the CFD software or wind tunnel, these methods are resource-intensive, time-consuming, and cannot cover all possible flight scenarios. Therefore, we propose to present the forces in the $x,y,z$ directions as three independent triangular functions with random periods and amplitudes, as shown in Table I. Since downwash flow mainly acts in the vertical direction, the force disturbance in the $z$ direction is more significant. For the torque disturbances around three axes, we simulate them using three independent Gaussian noises. This disturbance randomization gets rid of extensive randomization of aerodynamic and trajectory parameters of two quadcopter. Instead, it randomizes the external forces in the approach-hover-leave process numerically according to the conclusion in [6]. For example, variations in the approach and departure speeds of the upper quadcopter as well as the intensity of the downwash flow field affect the rate of change of external force disturbances. This is equivalent to randomizing the force disturbance period. Meanwhile, factors such as the relative hovering altitude and platform shape of the upper can determine the magnitude of the external force disturbances, we can therefore randomize the amplitude and maximum force directly. However, the shape of two drone and their relative horizontal position and yaw angle determine the torque disturbance significantly. Since our setup differs significantly from that in [6], we simulate them with Gaussian noise instead.
@@ III-A Simulation Experiments and Analysis
We conduct an experiment in simulator to verify that the proposed disturbances contribute to making the residual module adaptive to high-fidelity simulated disturbances. We approximate the effect of real downwash flow as a function of the relative positions and model parameters of two quadcopters. In our setup, the LQ takes off from the ground and hovers at $(0,0,1.2)$, while the SQ takes off vertically from $(-1,0,0)$ and flies along the positive $x$-axis at a speed of $0.2$ m/s at a certain height. After hovering above the LQ for 5 seconds, the SQ continues along the positive $x$-axis and lands vertically at $(1,0,0)$. We set the vertical distance between the SQ and LQ as $\left\{0.25,0.5,0.75\right\}$ and the relative yaw angle is always 00. Under this setup, the disturbance will be quite significant and traceable, so that we can directly apply the conclusions from [6] and reveal the corresponding compensatory behaviors generated by the residual module.
@@ III-B1 Comparative Experiments
Figure 5: The procedure of quadcopter mid-air docking. Stage 1: Taking off and hovering. The residual RL controller only assists the basic controller on the large quadcopter (LQ) to achieve faster responses. Stage 2: Small quadcopter approaching. Small quadcopter (SQ) hovers above the LQ and vertically approaches it, which generates strong downwash flow disturbances, and the residual RL controller generates thrust and rate compensation to help stabilize LQ’s position and attitude. (LPF: low-pass filtered) Stage 3: Docked with LQ. The SQ falls freely from 5cm above the LQ, gets docked with the LQ and generates an impulse. The overall mass changes and the thrust compensation from the RL controller on LQ reaches the peak. Stage 4: SQ leaving. SQ takes off from the LQ vertically, and the downwash airflow reappears and gradually decreases while the thrust compensation from the RL controller also gradually decreases.
@@ IV Conclusion
This paper proposes the ProxFly, which leverages a cascaded controller combined with residual RL for diverse close proximity tasks, including hovering, trajectory tracking, in-air docking, and undocking. This approach does not rely on any communication or accurate knowledge of system parameters but only uses a lightweight MLP to learn how to compensate for high-level commands. ProxFly leverages the strengths of both classical control and deep RL. On the one hand, it requires less domain expertise and avoids time-consuming controller parameter tuning. On the other hand, compared to end-to-end RL, it improves sampling efficiency and reduces the proportion of unexplained signals from the black box in the overall control signals. The demonstration of this paper can be found at https://youtu.be/NhPKgzd3l6w.